GROUPS AS OF FRIDAY, MARCH 10

A ALPHA:  Ballenger, Price
B BETA:  Burford, McCoy
Γ GAMMA:  Perez, Williams
Δ DELTA:
E EPSILON:  Jones, Joseph
Z ZETA:  Blevins, Miller
H ETA:  Smith, Walicek
Θ THETA:  Marshall, Mazurowski
I IOTA:
K KAPPA:  Kruppa, Sutherland
Λ LAMBDA:  McWhorter, Skinner
M MU:  Malcom, Nadeau
N NU:  Goines, Nguyen, Williamson
Ξ XI:  Brady, Drew
O OMICRON: Bezner, Deeke, Harris
Π PI:  Hobson, Valderrama
P RHO:
Σ SIGMA:
T TAU:
Υ UPSILON:
Φ PHI:
X CHI:
Ψ PSI:
Ω OMEGA: